# On the Lebesgue Constant of Extended-Domain Spectral Methods for Elliptic PDEs

### v2

@@ 4 Pre-Asymptotic Stability of the Convection-Diffusion Operator
Lemma 4.1 (Decay of Inverse Entries [17, Prop. 2]).
Let $\{\mathbf{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\mathbf{M}_{N}$ satisfy
[…]
@@
$\sup_{N\in\mathbb{N}}\left\lVert\mathbf{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
@@
The machinery of the proof follows [17]. Please see A. ∎
To apply this lemma, we must demonstrate that the sequence of physical-space operators, $\{\mathbf{L}_{CD,N}\}_{N\in\mathbb{N}}$, satisfies both of its hypotheses. These properties rely on foundational results for our specific spectral scheme, which we establish first.
@@
The space $V_{N}$, after a linear change of variables from $I=(0,L)$ to $0,2\pi]$, is a space of trigonometric polynomials of degree at most $N$. For such spaces, the magnitude of the cardinal functions can be related to the Green’s function with poles at the interpolation nodes. A foundational result in potential theory (see, e.g., [18, Chap. 4]) establishes an asymptotic relationship. For a cardinal function $p_{j}(x)$ corresponding to a set of quasi-uniform nodes $\{x_{k}\}$, this relationship is given by:
$\log|p_{j}(x)|=(N-1)\left(U^{\nu_{j}}(x_{j})-U^{\nu_{j}}(x)\right)+\epsilon_{N}(x,x_{j}),$ (6)
@@
$g(x)-g(x_{j})\geq c_{0}(x-x_{j})^{2},\quad\forall x,x_{j}\in I^{\prime}.$ (7)
@@
Substituting this into the relation (6) and using the convexity bound (7), we get:
@@
The exponential term in $N$ will always overwhelm the polynomial prefactor $N^{2C_{2}}$. This proves strong, Gaussian-like exponential decay away from the center point $x_{j}$. A simpler linear exponential decay bound, as stated in the lemma, follows directly from this stronger result for $d$ in a bounded domain. ∎
@@
The integrand $g$ is a trigonometric polynomial of degree at most $2N$, and the error $E_{N}(g)$ arises from aliasing. The key to bounding this error is to note that it can be controlled by the norms of the functions being multiplied. A standard result from the analysis of pseudospectral methods, derived by decomposing the solution into low- and high-frequency components, provides the following bound (see the argument in [16, Sec. 6.4]). For any $\epsilon>0$, there exists a constant $C_{\epsilon}$, independent of $N$, such that:
@@
We now show $u$ is a weak solution to $\mathcal{L}_{CD}u=0$. For any fixed smooth test function $v\in C_{c}^{\infty}(I)$, weak convergence implies $a(u_{N},v)\to a(u,v)$. The quadrature error for fixed $v$ is negligible for large $N$, so $a(u_{N},v)\approx a_{Q}(u_{N},\mathcal{I}_{N}v)=(f_{I,N},\mathcal{I}_{N}v)_{Q}$. As $N\to\infty$, this last term vanishes because $\left\lVert f_{I,N}\right\rVert_{L^{2}}\to 0$ and $\left\lVert\mathcal{I}_{N}v\right\rVert_{L^{2}}$ is bounded. Thus, $a(u,v)=0$ for all $v$. This means $\mathcal{L}_{CD}u=0$ weakly. However, the continuous problem is well-posed, with $u=0$ as its only solution. This contradicts our finding that $\left\lVert u\right\rVert_{L^{2}}\neq 0$. The initial assumption of instability must be false. Therefore, the scheme is stable. ∎
@@
Let $\mathbf{L}_{N}=\mathbf{A}_{N}\mathbf{W}_{N}^{-1}$ be the physical-space representation of a linear differential operator $\mathcal{L}$ with analytic coefficients. The entries of $\mathbf{L}_{N}$ decay exponentially away from the main diagonal. That is, there exist constants $C_{L}>0$ and $\beta>0$, independent of $N$, such that
$|(\mathbf{L}_{N})_{ij}|\leq C_{L}e^{-\beta|x_{i}-x_{j}|}.$
@@
The entry of the physical-space operator is $(\mathbf{L}_{N})_{ij}=(\mathcal{L}p_{j})(x_{i})$, where $p_{j}(x)$ is the cardinal interpolant in $V_{N}$. In Lemma 4.2, we established that $p_{j}(x)$ decays exponentially away from its center point $x_{j}$. Since $\mathcal{L}$ is a local, finite-order differential operator, its application preserves this exponential localization. By Cauchy’s differentiation formula, the derivatives of an analytic function decay at the same rate as the function itself. Thus, $(\mathcal{L}p_{j})(x_{i})$ is exponentially small for $|x_{i}-x_{j}|\gg 0$, satisfying the first hypothesis of Lemma 4.1. ∎
@@
For the convection-diffusion operator $\mathcal{L}_{CD}=-d^{2}/dx^{2}+k(x)d/dx$ with a bounded and strictly positive convection coefficient $k(x)$, the sequence of inverse physical-space operators $\{\mathbf{L}_{CD,N}^{-1}\}$ is uniformly bounded in the induced $\ell_{\infty}$-norm. That is, there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\mathbf{L}_{CD,N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
@@
The proof relies on the stability of the numerical scheme established in Lemma 4.3. Let $\mathbf{u}=\mathbf{L}_{CD,N}^{-1}\mathbf{f}$, which corresponds to the solution $u_{N}\in V_{N}$ for an interpolated forcing function $f_{I}(x)$. From Lemma 4.3, we have the stability estimate
$\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}.$ (8)
The proof proceeds by connecting the discrete $\ell_{\infty}$-norm to the continuous norms. By the one-dimensional Sobolev embedding theorem, $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}$. The solution vector is bounded by $\left\lVert\mathbf{u}\right\rVert_{\infty}\leq\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}$. The forcing term is bounded via the quadrature rule by $\left\lVert f_{I}\right\rVert_{L^{2}(I)}^{2}=\sum f_{i}^{2}w_{i}\leq\left\lVert\mathbf{f}\right\rVert_{\infty}^{2}L$. Combining these inequalities yields:
$\left\lVert\mathbf{u}\right\rVert_{\infty}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{emb}C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}\leq(C_{emb}C_{num}\sqrt{L})\left\lVert\mathbf{f}\right\rVert_{\infty}.$
@@
$\left\lVert\mathbf{L}_{CD,N}^{-1}\right\rVert_{\infty}=\sup_{\mathbf{f}\neq 0}\frac{\left\lVert\mathbf{u}\right\rVert_{\infty}}{\left\lVert\mathbf{f}\right\rVert_{\infty}}\leq C_{emb}C_{num}\sqrt{L}.$
Since this bound is independent of $N$, the second hypothesis of Lemma 4.1 is satisfied. ∎
@@
Let $\mathbf{L}_{CD,N}=\mathbf{A}_{CD,N}\mathbf{W}_{N}^{-1}$ be the physical-space representation of the constant-coefficient convection-diffusion ($k\neq 0$) operator on an ordered grid of collocation points $\{x_{i}\}$. The inverse matrix $\mathbf{L}_{CD,N}^{-1}$, which is the discrete Green’s function, is numerically quasi-sparse. There exist constants $C>0$ and $\alpha>0$, independent of $N$, such that its entries are bounded by:
$|(\mathbf{L}_{CD,N}^{-1})_{ij}|\leq Ce^{-\alpha|x_{i}-x_{j}|}.$
@@
The hypotheses of Lemma 4.1 have been established for the sequence of matrices $\{\mathbf{L}_{CD,N}\}_{N\in\mathbb{N}}$. Specifically, Lemma 4.4 establishes the exponential decay of the matrix entries, and Lemma 4.5 establishes the uniform boundedness of the inverse. The conclusion of Lemma 4.1 therefore applies directly, which proves the theorem. ∎
@@
The structural dichotomy in the discrete Green’s functions explains the observed pre-asymptotic difference in stability. While both operators are asymptotically unstable, for any finite $N$, the physical-space operator for convection-diffusion, $\mathbf{L}_{CD}$, is significantly better conditioned than its Poisson counterpart, $\mathbf{L}_{P}$.
@@
For the discrete Green’s function of the convection-diffusion operator, $\mathbf{L}_{CD}^{-1}$, the exponential decay proven in Theorem 4.1 ensures that the row sums are bounded independently of $N$:
$\left\lVert\mathbf{L}_{CD}^{-1}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\mathbf{L}_{CD}^{-1})_{ij}|\leq\max_{i}\sum_{j=1}^{N}Ce^{-\alpha|x_{i}-x_{j}|}=O(1).$
In contrast, the inverse of the discrete Poisson operator, $\mathbf{L}_{P}^{-1}$, reflects its continuous counterpart’s global influence. Its entries exhibit slow algebraic decay, causing the row sums and thus $\left\lVert\mathbf{L}_{P}^{-1}\right\rVert_{\infty}$ to grow with $N$.
The condition number of the underlying collocation matrix is $\kappa(\mathbf{A})=\left\lVert\mathbf{A}\right\rVert\left\lVert\mathbf{A}^{-1}\right\rVert$. We can write $\mathbf{A}^{-1}=\mathbf{W}^{-1}\mathbf{L}^{-1}$, which gives the bound $\left\lVert\mathbf{A}^{-1}\right\rVert\leq\left\lVert\mathbf{W}^{-1}\right\rVert\left\lVert\mathbf{L}^{-1}\right\rVert$. The term $\left\lVert\mathbf{W}^{-1}\right\rVert$ is exponentially large and is the source of the ultimate asymptotic instability for both operators. However, in the pre-asymptotic regime, the stark difference between a bounded $\left\lVert\mathbf{L}_{CD}^{-1}\right\rVert_{\infty}$ and a growing $\left\lVert\mathbf{L}_{P}^{-1}\right\rVert_{\infty}$ explains the profoundly better practical stability of the convection-diffusion system. ∎
@@
5 Extension to Higher Dimensions
The one-dimensional analysis of the preceding sections extends to $d$-dimensional problems on tensor-product domains. The core findings—namely, the true asymptotic exponential instability of the method for both operators and the pre-asymptotic structural dichotomy of their inverses—are preserved. This extension relies on the separable structure of the basis, grid, and operators.
5.1 Framework in $d$-Dimensions

### v3

@@ 4. Pre-Asymptotic Stability of the Convection-Diffusion Operator
Lemma 4.1 (Decay of Inverse Entries [jaffard1990proprietes, Prop. 2]).
Let $\{\bm{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\bm{M}_{N}$ satisfy
$|(\bm{M}_{N})_{ij}|\leq C_{M}e^{-\beta|i-j|}\quad\text{for all }1\leq i,j\leq N.$
@@
$\sup_{N\in\mathbb{N}}\left\lVert\bm{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
@@
The machinery of the proof follows [jaffard1990proprietes]. Please see A. Q.E.D.
To apply this lemma, we must demonstrate that the sequence of physical-space operators, $\{\bm{L}_{CD,N}\}_{N\in\mathbb{N}}$, satisfies both of its hypotheses. These properties rely on foundational results for our specific spectral scheme, which we establish first.
@@
The space $V_{N}$, after a linear change of variables from $I=(0,L)$ to $0,2\pi]$, is a space of trigonometric polynomials of degree at most $N$. For such spaces, the magnitude of the cardinal functions can be related to the Green’s function with poles at the interpolation nodes. A foundational result in potential theory (see, e.g., [saff1997logarithmic, Chap. 4]) establishes an asymptotic relationship. For a cardinal function $p_{j}(x)$ corresponding to a set of quasi-uniform nodes $\{x_{k}\}$, this relationship is given by:
$\log|p_{j}(x)|=(N-1)\left(U^{\nu_{j}}(x_{j})-U^{\nu_{j}}(x)\right)+\epsilon_{N}(x,x_{j}),$ (4.1)
@@
$g(x)-g(x_{j})\geq c_{0}(x-x_{j})^{2},\quad\forall x,x_{j}\in I^{\prime}.$ (4.2)
@@
Substituting this into the relation (4.1) and using the convexity bound (4.2), we get:
@@
The exponential term in $N$ will always overwhelm the polynomial prefactor $N^{2C_{2}}$. This proves strong, Gaussian-like exponential decay away from the center point $x_{j}$. A simpler linear exponential decay bound, as stated in the lemma, follows directly from this stronger result for $d$ in a bounded domain. Q.E.D.
@@
The integrand $g$ is a trigonometric polynomial of degree at most $2N$, and the error $E_{N}(g)$ arises from aliasing. The key to bounding this error is to note that it can be controlled by the norms of the functions being multiplied. A standard result from the analysis of pseudospectral methods, derived by decomposing the solution into low- and high-frequency components, provides the following bound (see the argument in [canuto2006spectral, Sec. 6.4]). For any $\epsilon>0$, there exists a constant $C_{\epsilon}$, independent of $N$, such that:
@@
We now show $u$ is a weak solution to $\mathcal{L}_{CD}u=0$. For any fixed smooth test function $v\in C_{c}^{\infty}(I)$, weak convergence implies $a(u_{N},v)\to a(u,v)$. The quadrature error for fixed $v$ is negligible for large $N$, so $a(u_{N},v)\approx a_{Q}(u_{N},\mathcal{I}_{N}v)=(f_{I,N},\mathcal{I}_{N}v)_{Q}$. As $N\to\infty$, this last term vanishes because $\left\lVert f_{I,N}\right\rVert_{L^{2}}\to 0$ and $\left\lVert\mathcal{I}_{N}v\right\rVert_{L^{2}}$ is bounded. Thus, $a(u,v)=0$ for all $v$. This means $\mathcal{L}_{CD}u=0$ weakly. However, the continuous problem is well-posed, with $u=0$ as its only solution. This contradicts our finding that $\left\lVert u\right\rVert_{L^{2}}\neq 0$. The initial assumption of instability must be false. Therefore, the scheme is stable. Q.E.D.
@@
Let $\bm{L}_{N}=\bm{A}_{N}\bm{W}_{N}^{-1}$ be the physical-space representation of a linear differential operator $\mathcal{L}$ with analytic coefficients. The entries of $\bm{L}_{N}$ decay exponentially away from the main diagonal. That is, there exist constants $C_{L}>0$ and $\beta>0$, independent of $N$, such that
$|(\bm{L}_{N})_{ij}|\leq C_{L}e^{-\beta|x_{i}-x_{j}|}.$
@@
The entry of the physical-space operator is $(\bm{L}_{N})_{ij}=(\mathcal{L}p_{j})(x_{i})$, where $p_{j}(x)$ is the cardinal interpolant in $V_{N}$. In Lemma 4.2, we established that $p_{j}(x)$ decays exponentially away from its center point $x_{j}$. Since $\mathcal{L}$ is a local, finite-order differential operator, its application preserves this exponential localization. By Cauchy’s differentiation formula, the derivatives of an analytic function decay at the same rate as the function itself. Thus, $(\mathcal{L}p_{j})(x_{i})$ is exponentially small for $|x_{i}-x_{j}|\gg 0$, satisfying the first hypothesis of Lemma 4.1. Q.E.D.
@@
For the convection-diffusion operator $\mathcal{L}_{CD}=-d^{2}/dx^{2}+k(x)d/dx$ with a bounded and strictly positive convection coefficient $k(x)$, the sequence of inverse physical-space operators $\{\bm{L}_{CD,N}^{-1}\}$ is uniformly bounded in the induced $\ell_{\infty}$-norm. That is, there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\bm{L}_{CD,N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
@@
The proof relies on the stability of the numerical scheme established in Lemma 4.3. Let $\bm{u}=\bm{L}_{CD,N}^{-1}\bm{f}$, which corresponds to the solution $u_{N}\in V_{N}$ for an interpolated forcing function $f_{I}(x)$. From Lemma 4.3, we have the stability estimate
$\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}.$ (4.3)
The proof proceeds by connecting the discrete $\ell_{\infty}$-norm to the continuous norms. By the one-dimensional Sobolev embedding theorem, $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}$. The solution vector is bounded by $\left\lVert\bm{u}\right\rVert_{\infty}\leq\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}$. The forcing term is bounded via the quadrature rule by $\left\lVert f_{I}\right\rVert_{L^{2}(I)}^{2}=\sum f_{i}^{2}w_{i}\leq\left\lVert\bm{f}\right\rVert_{\infty}^{2}L$. Combining these inequalities yields:
$\left\lVert\bm{u}\right\rVert_{\infty}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{emb}C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}\leq(C_{emb}C_{num}\sqrt{L})\left\lVert\bm{f}\right\rVert_{\infty}.$
@@
$\left\lVert\bm{L}_{CD,N}^{-1}\right\rVert_{\infty}=\sup_{\bm{f}\neq 0}\frac{\left\lVert\bm{u}\right\rVert_{\infty}}{\left\lVert\bm{f}\right\rVert_{\infty}}\leq C_{emb}C_{num}\sqrt{L}.$
Since this bound is independent of $N$, the second hypothesis of Lemma 4.1 is satisfied. Q.E.D.
@@
Let $\bm{L}_{CD,N}=\bm{A}_{CD,N}\bm{W}_{N}^{-1}$ be the physical-space representation of the constant-coefficient convection-diffusion ($k\neq 0$) operator on an ordered grid of collocation points $\{x_{i}\}$. The inverse matrix $\bm{L}_{CD,N}^{-1}$, which is the discrete Green’s function, is numerically quasi-sparse. There exist constants $C>0$ and $\alpha>0$, independent of $N$, such that its entries are bounded by:
$|(\bm{L}_{CD,N}^{-1})_{ij}|\leq Ce^{-\alpha|x_{i}-x_{j}|}.$
@@
The hypotheses of Lemma 4.1 have been established for the sequence of matrices $\{\bm{L}_{CD,N}\}_{N\in\mathbb{N}}$. Specifically, Lemma 4.4 establishes the exponential decay of the matrix entries, and Lemma 4.5 establishes the uniform boundedness of the inverse. The conclusion of Lemma 4.1 therefore applies directly, which proves the theorem. Q.E.D.
Figure 2. Direct visualization of the structural dichotomy in the physical-space operators and their inverses for $N=64$ and $\delta=2.0$. The colormaps show $\log_{10}$ of the absolute value of the matrix entries. (a), (c): Both forward operators are localized. (b): The inverse of the Poisson operator is dense. (d): The inverse of the convection-diffusion operator is numerically quasi-sparse, visually confirming the exponential decay predicted in Theorem 4.6.
Remark 4.7 (Visual Verification of the Green’s Function Structure).
The structural dichotomy established by Theorem 4.6 is strikingly visible in numerical computations. Figure 2 displays the magnitude of the entries for the physical-space operators and their inverses.
  • Poisson Operator (Fig. 2b): The inverse matrix $(\bm{L}_{P,N})^{-1}$ is dense with slow algebraic decay away from the diagonal, reflecting the global nature of the elliptic Green’s function.
  • Convection-Diffusion Operator (Fig. 2d): In contrast, the inverse matrix $(\bm{L}_{CD,N})^{-1}$ is numerically quasi-sparse. The entries exhibit the sharp exponential decay predicted by our analysis.
This sparsity in the inverse operator is the physical mechanism that suppresses the accumulation of round-off errors and stabilizes the method in the pre-asymptotic regime, distinct from the behavior of the underlying ill-conditioned basis.
@@
The structural dichotomy in the discrete Green’s functions explains the observed pre-asymptotic difference in stability. While both operators are asymptotically unstable, for any finite $N$, the physical-space operator for convection-diffusion, $\bm{L}_{CD}$, is significantly better conditioned than its Poisson counterpart, $\bm{L}_{P}$.
@@
For the discrete Green’s function of the convection-diffusion operator, $\bm{L}_{CD}^{-1}$, the exponential decay proven in Theorem 4.1 ensures that the row sums are bounded independently of $N$:
$\left\lVert\bm{L}_{CD}^{-1}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\bm{L}_{CD}^{-1})_{ij}|\leq\max_{i}\sum_{j=1}^{N}Ce^{-\alpha|x_{i}-x_{j}|}=O(1).$
In contrast, the inverse of the discrete Poisson operator, $\bm{L}_{P}^{-1}$, reflects its continuous counterpart’s global influence. Its entries exhibit slow algebraic decay, causing the row sums and thus $\left\lVert\bm{L}_{P}^{-1}\right\rVert_{\infty}$ to grow with $N$.
The condition number of the underlying collocation matrix is $\kappa(\bm{A})=\left\lVert\bm{A}\right\rVert\left\lVert\bm{A}^{-1}\right\rVert$. We can write $\bm{A}^{-1}=\bm{W}^{-1}\bm{L}^{-1}$, which gives the bound $\left\lVert\bm{A}^{-1}\right\rVert\leq\left\lVert\bm{W}^{-1}\right\rVert\left\lVert\bm{L}^{-1}\right\rVert$. The term $\left\lVert\bm{W}^{-1}\right\rVert$ is exponentially large and is the source of the ultimate asymptotic instability for both operators. However, in the pre-asymptotic regime, the stark difference between a bounded $\left\lVert\bm{L}_{CD}^{-1}\right\rVert_{\infty}$ and a growing $\left\lVert\bm{L}_{P}^{-1}\right\rVert_{\infty}$ explains the profoundly better practical stability of the convection-diffusion system. Q.E.D.
Remark 4.9 (Practical Implications for Stabilization).
This structural dichotomy has direct practical consequences for solver design. Because the quasi-sparse inverse of the convection-diffusion operator effectively filters the high-frequency noise inherent to the ill-conditioned Fourier extension frame, this operator requires significantly less oversampling or regularization parameter tuning than the Poisson operator to achieve stable, high-accuracy solutions in the pre-asymptotic regime.
@@
5. Extension to Higher Dimensions
The structural dichotomy observed in one dimension extends to $d$-dimensional problems on tensor-product domains. The analysis here reveals a tension: while the baseline ill-conditioning of the linear system becomes even more severe due to the tensor-product nature of the basis, the favorable pre-asymptotic properties of the convection-diffusion operator are preserved due to the discrete maximum principle.
5.1. Framework in $d$-Dimensions